OpenStudy (anonymous):

lim (x^2-4)/(x^3+8) x->-2

OpenStudy (anonymous):

i wrote dnehere too????

myininaya (myininaya):

{(x-2)(x+2)}/{(x+2)(x^2-2x+4)} x+2 cancel so we have (x-2)/(x^2-2x+4) -4/(4+4+4) -4/12 -1/3

OpenStudy (anonymous):

marked wrong!!

myininaya (myininaya):

azkhan the last problem had 90 in it right? i didnt understand your response

OpenStudy (anonymous):

isn that the same thing we had todo with the other questin????

OpenStudy (anonymous):

noooo it had 9

OpenStudy (anonymous):

sorry typo!!!

myininaya (myininaya):

oh you wrote 90

myininaya (myininaya):

that makes a difference the limit does exist then

myininaya (myininaya):

(t^2-9)/(t-3)={(t-3)(t+3)}/{t-3}=t+3 as t->3, f(t)->3+3=6 limit is 6

OpenStudy (anonymous):

(a^3+b^3)=(a-b) ...........??????

OpenStudy (anonymous):

(a^2-2a+b^2)??????

OpenStudy (anonymous):

Sum of Cubes: (a³ + b³) = (a + b)(a² - ab + b²).

OpenStudy (anonymous):

thank u!!!!

OpenStudy (anonymous):

use l'hospital expression bcomes 2x/3x^2 =2/3x applying the limit, 2/3*-2=-1/3